OpenStudy (anonymous):

Multiply and simplify: (4y − 1)(8y + 3) (Points : 1) 32y2 − 4y − 3 32y2 + 4y + 3 32y2 − 4y + 3 32y2 + 4y − 3

OpenStudy (anonymous):

(4y-1)(8y+3)=4y(8y)+4y(3)+(-1)(8y)+(-1)(3)=32y^2 +12y - 8y -3=32y^2 +4y -3